# Isomorphism classes of generating sets

### v1

@@ 2. Some preliminaries
[…]
For $(2)$, let $g:\mathbb{D}^{\prime}\to\mathbb{D}$ be an embedding. Let $A$ be the closure of $\operatorname{rng}(g)$ under join and meets. That is, $A$ is the increasing union of sets $A_{n}$, such that $A_{0}=\operatorname{rng}(g)$, and $A_{n+1}=\{a\vee b\mid a,b\in A_{n}\}\cup\{a\wedge b\mid a,b\in A\}$. Clearly, $A$ is closed under meets and joins. We claim that $\operatorname{rng}(g)$ is cofinal in $A$. For this, it suffices to prove that for every $n$, $A_{n}$ is cofinal in $A_{n+1}$. Indeed, since $D^{\prime}$ is directed, also $\operatorname{rng}(g)$ is directed and therefore $A_{0}$ is cofinal in $A_{1}$, and therefore $A_{1}$ stays directed. Inductively assume that $\operatorname{rng}(g)$ is cofinal in $A_{n}$, then by directness, it is cofinal in $A_{n+1}$.
By $(1)$, $A$ is a degree $2$ lattice. Hence $\mathbb{D}^{\prime}$ can be embedded cofinally into a degree $2$ lattice. ∎
Example 2.9.
Consider $\mathbb{D}=(\bigcup_{\alpha<\lambda_{0}}\{\alpha\}\times\omega_{\alpha},\leq_{D})$, where $\leq_{D}$, as in the previous example is the pointwise order on pairs of ordinals. Note that for every $(\alpha,\beta),(\gamma,\delta)\in D$, $(\max\{\alpha,\gamma\},\max\{\beta,\delta\}),(\min\{\alpha,\gamma\},\min\{\beta,\delta\})\in D$. Hence it is evident (and follows Proposition 2.8(2)) that the degree $2$ property of $\lambda_{0}\times(\sup_{\alpha<\lambda_{0}}\omega_{\alpha})$ is inherited to $\mathbb{D}$.
Corollary 2.10.
Suppose that $\mathbb{D}$ is a two-dimensional, directed, well-founded poset, then $\mathbb{D}$ embeds cofinally into a degree 2 sub-lattice of a product of cardinals.
Proof.
[…]
$\gamma=\sup\{\alpha+1\mid\exists(a^{\prime},b^{\prime})\in D_{<(a,b)},\ a^{\prime}<_{A}a,\exists\beta,g(a^{\prime},b^{\prime})=(\alpha,\beta)\},$
$\delta=\sup\{\beta+1\mid\exists(a^{\prime},b^{\prime})\in D_{<(a,b)},\ b^{\prime}<_{B}b,\exists\alpha,g(a^{\prime},b^{\prime})=(\alpha,\beta)\}.$
3. Finite support D-based iteration
In this section we will describe a forcing iteration which is indexed by a non-linear partial order. As we denoted in the introduction, similar iterations appeared in various places in the literature. Notably, in the works of Shelah (see, for example, [41, 42]), this type of iteration is modeled as linear iteration with restricted memory — so in Shelah’s language, asking the memory to be transitive is parallel to forcing along a partial order. In [42], Shelah discribed a more general mechanism that allows one to deal with certain schemes which might be ill–founded or non-transitive.
[…]
  • $\mathbb{P}_{<d}$ is the set of finitely supported sequences $p=\langle p_{e}\mid e\in D_{<d}\rangle$ such that for all $e\in D_{<d}$, $\Vdash_{\mathbb{P}_{<e}}p(e)\in\dot{\mathbb{Q}}_{e}$.
  • $1_{\mathbb{P}_{<d}}\Vdash\dot{\mathbb{Q}}_{d}\text{ is a partial order}$.
We denote by $\mathbb{P}_{\mathbb{D}}$ the set of finitely supported sequences $p=\langle p_{e}\mid e\in D\rangle$ such that for all $e\in D$, $p\restriction D_{<e}\Vdash_{\mathbb{P}_{<e}}p(e)\in\dot{\mathbb{Q}}_{e}$. We will also write $\mathbb{P}_{\leq d}$ instead of $\mathbb{P}_{<d}*\dot{\mathbb{Q}}_{d}$. The order is defined by $p\leq_{\mathbb{P}_{\mathbb{D}}}q$ iff for every $d\in D$, $p\restriction D_{<d}\Vdash_{\mathbb{P}_{<d}}p(d)\leq_{\dot{\mathbb{Q}_{d}}}q(d)$.
Ultimately, we will only consider iterations based on well-founded partially ordered sets, as the names ˙$\dot{\mathbb{Q}}_{d}$ will be defined by well-founded recursion.
Lemma 3.1.
Let $\mathbb{D}^{\prime}\subseteq\mathbb{D}$ be downwards closed. Then $P_{\mathbb{D}}$ projects via restriction to $\mathbb{P}_{\mathbb{D}^{\prime}}.$
Proof.
We will check that the restriction map from $\mathbb{P}_{\mathbb{D}}$ to $\mathbb{P}_{\mathbb{D}^{\prime}}$ is a projection. Clearly, it is order preserving. Suppose that $q\in\mathbb{P}_{D^{\prime}}$ is stronger than $p\restriction D^{\prime}$. We will find an extension $p^{\prime}\leq p$ in $\mathbb{P}_{\mathbb{D}}$ such that $p^{\prime}\restriction D^{\prime}=q$.
Define $p^{\prime}=q\cup p\restriction(D\setminus D^{\prime})$. Then by definition, $p^{\prime}$ belongs to $\mathbb{P}_{\mathbb{D}}$ Also $p^{\prime}\restriction D^{\prime}=q$. It remains to prove that for every $d\in D$,
@@ 3. Finite support D-based iteration
This forcing is $\kappa$-closed, and $\kappa^{+}$-cc. It adds a set $X$ which is a $\subseteq^{*}$-lower bound for the filter $F$.
@@
  1. I.
    $\dot{U}_{<d}$ is an $\mathbb{P}_{<d}$-name for an ultrafilter on $\omega$ that extends $\bigcup_{e<_{D}d}\dot{U}_{e}$.
  2. II.
    ˙$\dot{\mathbb{Q}}_{d}$ is a $\mathbb{P}_{<d}$-name for the Mathias forcing $\mathbb{M}(\dot{U}_{<d})$.
  3. III.
    $\dot{U}_{d}$ is an $\mathbb{P}_{\leq d}$-name for an ultrafilter extending $\dot{U}_{<d}\cup\{\dot{B}_{d}\}$ where $\dot{B}_{d}=\{(p,\check{\alpha})\mid p\restriction D_{<d}\Vdash``\check{\alpha}\text{ appears in the stem of }p(d)"\}$ is the canonical name for an ˙$\dot{\mathbb{Q}}_{d}$-generic set.
Given $\mathbb{D}$ and a suitable sequence $\langle\dot{U}_{<d},\dot{U}_{d}:d\in D\rangle$, the associated iteration of Mathias forcings based on $\mathbb{D}$ is the iteration $\langle\mathbb{P}_{<d},\dot{\mathbb{Q}}_{d}:d\in D\rangle$ which is uniquely determined by the choice of $\langle\dot{U}_{<d},\dot{U}_{d}:d\in D\rangle$.
Fix such a suitable sequence $\langle\dot{U}_{<d},\dot{U}_{d}:d\in D\rangle$ and denote by $\mathbb{M}_{D}$ the resulting partial order $\mathbb{P}_{\mathbb{D}}$.
Example 3.3.
Let $\lambda_{0}$ and $\lambda_{1}$ be regular uncountable cardinals. Let $(D,{\leq_{D}})=(\lambda_{0},{\leq})\times(\lambda_{1},{\leq}),$ so $\mathbb{D}$ is the directed partial order on $\lambda_{0}\times\lambda_{1}$ given by
$(\alpha_{0},\alpha_{1})\leq_{D}(\beta_{0},\beta_{1})\;\Longleftrightarrow\;\alpha_{0}\leq\beta_{0}\,\wedge\,\alpha_{1}\leq\beta_{1}.$
We will eventually prove (Corollary 3.8) that for such $\mathbb{D}$, it is possible to construct a suitable sequence ${\langle}\dot{U}_{<d},\dot{U}_{d}\mid d\in D{\rangle}$.
${{(V^{\mathbb{P}_{2,0}},U_{2,0})}}$${{(V^{\mathbb{P}_{2,1}},U_{2,1})}}$${{(V^{\mathbb{P}_{2,2}},U_{2,2})}}$${{(V^{\mathbb{P}_{<(2,1)}},U_{<(2,1)})}}$${{(V^{\mathbb{P}_{<(2,2)}},U_{<(2,2)})}}$${{(V^{\mathbb{P}_{1,0}},U_{1,0})}}$${{(V^{\mathbb{P}_{1,1}},U_{1,1})}}$${{(V^{\mathbb{P}_{1,2}},U_{1,2})}}$${{(V^{\mathbb{P}_{<(1,1)}},U_{<(1,1)})}}$${{(V^{\mathbb{P}_{<(1,2)}},U_{<(1,2)})}}$${{(V^{\mathbb{P}_{0,0}},U_{0,0})}}$${{(V^{\mathbb{P}_{0,1}},U_{0,1})}}$${{(V^{\mathbb{P}_{0,2}},U_{0,2})}}$${{(V,U_{<(0,0)})}}$$\scriptstyle{\mathbb{M}_{U_{<(2,1)}}}$$\scriptstyle{\mathbb{M}_{U_{<(2,2)}}}$$\scriptstyle{\mathbb{M}_{U_{<(2,0)}}}$$\scriptstyle{\mathbb{M}_{U_{<(1,1)}}}$$\scriptstyle{\mathbb{M}_{U_{<(1,2)}}}$$\scriptstyle{\mathbb{M}_{U_{<(1,0)}}}$$\scriptstyle{\mathbb{M}_{U_{<(0,1)}}}$$\scriptstyle{\mathbb{M}_{U_{<(0,2)}}}$$\scriptstyle{\mathbb{M}_{U_{<(0,0)}}}$
Figure 1. Models and ultrafilters of the $\lambda_{0}\times\lambda_{1}$-based iteration

### v2

@@ 3. $\mathbb{D}$-based iteration
We will check that the restriction map from $\mathbb{P}_{\mathbb{D}}$ to $\mathbb{P}_{\mathbb{D}^{\prime}}$ is a projection. Clearly, it is order preserving. Suppose that $q\in\mathbb{P}_{\mathbb{D}^{\prime}}$ is stronger than $p\restriction D^{\prime}$. We will find an extension $p^{\prime}\leq p$ in $\mathbb{P}_{\mathbb{D}}$ such that $p^{\prime}\restriction D^{\prime}=q$.
Define $p^{\prime}=q\cup p\restriction(D\setminus D^{\prime})$. Then by definition, $p^{\prime}$ belongs to $\mathbb{P}_{\mathbb{D}}$, and $p^{\prime}\restriction D^{\prime}=q$. It remains to prove that for every $d\in D$,
@@
As one can see from the proof, this projection is a forking projection [40], that is to say given a condition $p$ and a condition $r\leq p\restriction D^{\prime}$, there is a weakest condition $q=r\cup(p\restriction D\setminus D^{\prime})$ which extends $p$ and project to $r$.
@@
This forcing is $\kappa$-closed, and $(\kappa^{<\kappa})^{+}$-cc. It adds a set $X$ which is a $\subseteq^{*}$-lower bound for the filter $F$. Note that $\mathbb{M}(F)$ is canonically $\kappa$-directed closed, namely, every directed set of conditions of size $\kappa$ has a greatest lower bound. Indeed, for a directed set of conditions $\langle s_{i},A_{i}\rangle$, their greatest lower bound is $\langle\bigcup s_{i},\bigcap A_{i}\rangle$.
Lemma 3.3.
@@
  1. (1)
    $\mathbb{P}_{\mathbb{D}}$ has a dense subset of $\mathbb{P}_{\mathbb{D}}$ of conditions $p$ such that for every $d\in\operatorname{dom}p$, $p(d)=\langle\check{s},\dot{A}\rangle$ (namely, the stem is a canonical name).
  2. (2)
    $\mathbb{P}_{\mathbb{D}}$ is canonically $\kappa$-directed closed.
  3. (3)
    $\mathbb{P}_{\mathbb{D}}$ is $(\kappa^{<\kappa})^{+}$-Knaster and in particular $(\kappa^{<\kappa})^{+}$-$c$.
Proof.
Let us prove item $(1)$ by induction on $\mathbb{D}_{+}=\mathbb{D}\cup\{\infty\}$, where $\infty$ is defined to be above all elements of $\mathbb{D}$.
Let $d\in\mathbb{D}_{+}$ and let us assume that for every $e<d$ the lemma holds for $\mathbb{P}_{<e}$. Let $p\in\mathbb{P}_{<d}$, and let $s=\operatorname{Supp}p=\langle e_{i}\mid i<i_{*}\rangle$, $i_{*}<\kappa$. Let us assume that we pick the enumeration so that $e_{i}<_{\mathbb{D}}e_{j}$ implies $i<j$.
Define a sequence of conditions $\langle p_{i}\mid i<i_{*}+1\rangle$ in $\mathbb{P}_{<d}$. Let $p_{0}=p$ and for every $i$, let $p_{i+1}\leq p_{i}$ such that
  • $p_{i+1}\restriction\mathbb{D}_{<e_{i}}$ decides the stem of $p_{i}(e_{i})$,
  • for every $e^{\prime}\in\operatorname{dom}p_{i+1}\cap\mathbb{D}_{<e_{i}}$, the stem of $p_{i+1}(e^{\prime})$ is a canonical name for a ground model set, and
  • $p_{i+1}\restriction(\mathbb{D}_{<d}\setminus\mathbb{D}_{<e_{i}})=p_{i}\restriction(\mathbb{D}_{<d}\setminus\mathbb{D}_{<e_{i}}).$
The second requirement is possible by the inductive hypothesis.
For a non-zero limit ordinal $i\leq i_{*}$, let $\operatorname{Supp}p_{i}=\bigcup_{j<i}\operatorname{Supp}p_{j}$ and for every $e\in\operatorname{Supp}p_{i}$, let $p_{i}(e)$ be the greatest lower bound of $\langle p_{j}(e)\mid j<i,\,e\in\operatorname{dom}p_{j}\rangle$. By the above remark, we can take the stem of $p_{i}(e)$ to be the canonical name for the union of the stems.
It is straightforward to verify that $p_{i}$ is indeed a condition which is stronger than $p_{j}$ for all $j<i$. Note that at step $i$, for every $e\in\operatorname{Supp}p_{i}$ such that $e\neq e_{k}$ for $k\geq i$, the stem of $p_{i}(e)$ is a canonical name for a ground model set, as wanted.
For $(2)$, note that for every $\in\mathbb{D}, ˙d\dot{\mathbb{Q}}_{d}$ is forced by $\mathbb{P}_{<d}$ to be canonically $\kappa$-directed closed and so a canonical lower bound can be defined coordinate-wise.
Let us prove $(3)$. Let $\mathcal{A}\subseteq\mathbb{P}_{\mathbb{D}}$ be a collection of conditions of cardinality $(\kappa^{<\kappa})^{+}$. Using the first part of the lemma, we may assume that for every $p\in\mathcal{A}$, and $e\in\operatorname{Supp}p$, the stem of $p(e)$ is a canonical name.
Since the support of the the conditions is of size $\kappa$, there is $\mathcal{S}\subseteq\mathcal{A}$ of cardinality $(\kappa^{<\kappa})^{+}$ such that $\langle\operatorname{Supp}p\mid p\in\mathcal{S}\rangle$ form a $\Delta$-system. Let $R$ be the root of the system. By shrinking $\mathcal{S}$, we may assume that there is a function $s\colon R\to[\kappa]^{<\kappa}$ such that for every $p\in\mathcal{S}$, and $e\in R$, $p(e)=\langle\check{s(e)},\dot{A}^{p}_{e}\rangle$. We claim that every pair of conditions $p,q\in\mathcal{S}$ is compatible. To see this let
$r(e)=\begin{cases}p(e)&e\in\operatorname{Supp}p\setminus R\\ q(e)&e\in\operatorname{Supp}q\setminus R\\ \langle\check{s(e)},\dot{A}^{p}_{e}\cap\dot{A}^{q}_{e}\rangle&e\in R.\end{cases}$
As each $\dot{F}_{e}$ is forced to be a filter, this is a condition stronger than $p$ and $q$. The same argument shows that any subset of $\mathcal{S}$ of cardinality less than $\kappa$ has a lower bound. ∎
As we will see later, the main challenge is to pick the filters $\dot{F}_{e}$ in a way that will allow us to isolate in the generic extension an ultrafilter which is generated by the Mathias generic sets.
Given an iteration $\mathbb{P}_{\mathbb{D}}$ of Mathias forcing, a generic filter $G\subseteq\mathbb{P}_{\mathbb{D}}$, and $d\in\mathbb{D}$, let
$G_{<d}=\{p\restriction D_{<d}\mid p\in G\}.$
By Lemma 3.1, it is a generic filter for $\mathbb{P}_{<d}$. Let $G_{d}=\{p(d)^{G_{<d}}\mid p\in G\}\subseteq\mathbb{Q}_{d}$, which is generic over $V[G_{<d}]$. Let $B_{d}=\bigcup\{s\mid\langle s,A\rangle\in G_{d}\}$ and fix a canonical name $\dot{B}_{d}$ for $B_{d}$.
Lemma 3.4.
Let $\mathbb{P}_{\mathbb{D}}$ be a $\kappa-$support iteration based on $\mathbb{D}$ of Mathias forcings relative to filters and let $G\subseteq\mathbb{P}_{\mathbb{D}}$ be a generic filter. In $V[G]$, for all $e,d\in\mathbb{D}$ such that $e\not\leq d$, $B_{d}\nsubseteq^{*}B_{e}$.
Proof.
Assume otherwise and let $p\in\mathbb{P}_{\mathbb{D}}$ force that $\dot{B}_{d}\setminus\dot{B}_{e}$ is bounded by some $\alpha<\kappa$. Let us look at $p(d)=\langle\check{s}_{d},\dot{A}_{d}\rangle$ and $p(e)=\langle\check{s}_{e},\dot{A}_{e}\rangle$. As $\dot{A}_{d}$ is a $\mathbb{P}_{<d}$-name for an unbounded subset of $\kappa$, there is an extension $r\leq p\restriction\mathbb{D}_{<d}$ in $\mathbb{P}_{<d}$ such that for some $\gamma>\alpha,\sup(s_{e})$, $r\Vdash\gamma\in\dot{A}_{d}$.
Let $\dot{A}^{\prime}_{e}$ be a $\mathbb{P}_{<e}$-name forced to be $\dot{A}_{e}\setminus\check{\gamma+1}$ and let $s^{\prime}_{d}=s_{d}\cup\{\gamma\}$. Let $q\leq p$ be the following condition:
$q(c)=\begin{cases}r(c)&c<d\\ \langle\check{s}^{\prime}_{d},\dot{A}_{d}\rangle&c=d\\ \langle\check{s}_{e},\dot{A}^{\prime}_{e}\rangle&c=e\\ p(c)&else.\end{cases}$
Then, $q\leq p$ and $q\Vdash\check{\gamma}\in\dot{B}_{d}\setminus\dot{B}_{e}$, contradiction. ∎
Corollary 3.5.
Let $\mathbb{P}_{\mathbb{D}}$ be as above and let $\mathbb{E}\subseteq\mathbb{D}$ such that
($\dagger$) $\text{for every }e_{0},e_{1}\in\mathbb{E},\text{ if }e_{0}<_{E}e_{1}\text{ then }\Vdash_{\mathbb{P}_{<e_{1}}}\dot{B}_{e_{0}}\in\dot{F}_{e_{1}}.$
Then, in the generic extension, $(\langle B_{e}\mid e\in\mathbb{E}\rangle,\supset^{*}\rangle\cong\mathbb{E}$.
Proof.
Let $e_{0},e_{1}\in\mathbb{E}$. If $e_{0}<_{E}e_{1}$ then by ($\dagger$3.5), and properties of $\mathbb{M}(\dot{F}_{e_{1}})$, $B_{e_{0}}\supset^{*}B_{e_{1}}$. Otherwise, by Lemma 3.4, $B_{e_{1}}\not\subseteq^{*}B_{e_{0}}$. ∎
Thus, the main task which is left is to show how the names $\dot{F}_{d}$ can be chosen so that the requirements of the corollary hold for a prescribed $\mathbb{E}$ and simultaneously generate an ultrafilter. In the next section we will formally define the iteration by stating precisely how the filters $\dot{F}_{d}$ are chosen.
4. Mathias iteration
The goal of this section is to define for a given well-founded, $\kappa^{+}$-directed partial order $\mathbb{D}$ a forcing notion that introduces an ultrafilter with a basis isomorphic to $\mathbb{D}$. We will be able to show that this forcing works for the cases $\kappa=\omega$ and $\kappa$ is indestructible supercompact.
Let $\mathbb{D}$ be $\kappa^{+}$-directed well-founded partial order. Find an ordinal $\mu$ with $|\mu|=|\mathbb{D}|$ and a linearization ${\langle}d_{\alpha}\mid\alpha<\mu{\rangle}$ of $\mathbb{D}$. Namely, ${\langle}d_{\alpha}\mid\alpha<\mu{\rangle}$ enumerates $\mathbb{D}$ and
($\triangle$) $\forall i,j<\mu,\ d_{i}\leq_{\mathbb{D}}d_{j}\Rightarrow i\leq j.$
In this situation we say that $\mu$ is a possible order type of $\mathbb{D}$. Note that by the $\kappa^{+}$-directedness if $\mu$ is a possible order type of $\mathbb{D}$, ${\rm cf}(\mu)>\kappa$. Let $\mathbb{D}^{+}=\mu^{+}\times\mathbb{D}$ with the usual product strict ordering, namely $(\alpha,d)<(\beta,e)$ if and only if $\alpha<\beta$ and $d<e$. Note that << is not the strict part of the product partial ordering $(\mu^{+},\leq)\times(\mathbb{D},\leq_{\mathbb{D}})$.
Let us assume further that the diamond principle $\diamondsuit_{\mu^{+}}(\mu^{+}\cap\operatorname{cof}(\mu))$ holds in $V$. Here $\operatorname{cof}(\mu)$ denotes the class of ordinals $\alpha$ with ${\rm cf}(\alpha)={\rm cf}(\mu)$. This form of diamond principle does not follow in general from the assumption that $2^{\mu}=\mu^{+}$, and in the sequel we will arrange it by forcing. We will define a non-linear iteration $\mathbb{P}=\mathbb{P}_{\mathbb{D}^{+}}$ of filter Mathias forcing indexed by $\mathbb{D}^{+}$, so by Lemma 3.3 $\mathbb{P}$ is $\kappa$-directed closed and $\kappa^{+}$-$c$.
We will use a form of diamond which is adapted to guessing subsets of $H_{\mu^{+}}$, which is possible because $|H_{\mu^{+}}|=\mu^{+}$. We will assume without loss of generality that $\mathbb{D}\in H_{\mu^{+}}$. It follows from $\diamondsuit_{\mu^{+}}(\mu^{+}\cap\operatorname{cof}(\mu))$ there are sequences $(A_{\alpha})_{\alpha\in\mu^{+}\cap\operatorname{cof}(\mu)}$ and $(M_{\alpha})_{\alpha<\mu^{+}}$ such that
  1. (1)
    $M_{\alpha}\prec H_{\mu^{+}}$ with $|M_{\alpha}|=\mu$ and $M_{\alpha}\cap\mu^{+}\in\mu^{+}$.
  2. (2)
    $(M_{\alpha})_{\alpha<\mu^{+}}$ is continuous and $\bigcup_{\alpha<\mu^{+}}M_{\alpha}=H_{\mu^{+}}$.
  3. (3)
    $\mathbb{D}\in M_{0}$ and $\mathbb{D}\cup\mu+1\subseteq M_{0}$.
  4. (4)
    $A_{\alpha}\subseteq M_{\alpha}$, and for every $A\subseteq H_{\mu^{+}}$ there are stationarily many $\alpha\in\mu^{+}\cap\operatorname{cof}(\mu)$ such that $A\cap M_{\alpha}=A_{\alpha}$.
The choice of the sequence $(M_{\alpha})_{\alpha<\mu^{+}}$ is not important, since such sequences always exist and any two such sequences agree on a club subset of $\mu^{+}$. Also note that $M_{\alpha}\cap\mathbb{D}^{+}=(M_{\alpha}\cap\mu^{+})\times\mathbb{D}$
[…]
As usual, ${\mathbb{P}}_{<a}$ is defined by recursion on $a\in\mathbb{D}^{+}$, with a minor modification that we are going to describe. As discussed above, the main part of the definition is choosing a ${\mathbb{P}}_{<a}$-name $\dot{F}_{a}$ for a $\kappa$-complete uniform filter on $\kappa$ in $V^{\mathbb{P}_{<a}}$.
Definition 4.1.
$\dot{F}_{a}$ is defined in $V$ by induction on $\beta<\mu^{+}$, and an subsidiary induction on $a\in\{\beta\}\times\mathbb{D}$. Assume that $\mathbb{P}_{\beta\times\mathbb{D}}$ was defined.
If $\beta\in\mu^{+}\cap\operatorname{cof}(\mu)$ and $A_{\beta}$ is a canonical ${\mathbb{P}}_{\beta\times\mathbb{D}}$-name for a uniform $\kappa$-complete ultrafilter on $\kappa$ (abbreviated by $\beta$ is an active stage”), we set for every $a=(\beta,d)\in\{\beta\}\times\mathbb{D}$,
$\dot{F}_{a}:=\{(1,\dot{A})\colon\dot{A}\text{ is a canonical }$ ${\mathbb{P}}_{\beta\times\mathbb{D}_{<d}}\text{-name}$
$\text{ for a subset of $\kappa$ and }\Vdash^{V}_{{\mathbb{P}}_{\beta\times\mathbb{D}}}\dot{A}\in A_{\beta}\}.$
If $\beta$ is not an active stage let $\dot{F}_{a}$ name the tail filter on $\kappa$.
The forcing expression $\Vdash^{V}_{{\mathbb{P}}_{\beta\times\mathbb{D}}}\dot{A}\in A_{\beta}$ is reasonable because $\beta\times\mathbb{D}_{<d}$ is a downwards closed subset of $\beta\times\mathbb{D}$, so that $\dot{A}$ makes sense as a ${\mathbb{P}}_{\beta\times\mathbb{D}}$-name for a subset of $\kappa$. It is easy to verify that in either case $\dot{F}_{a}$ is a ${\mathbb{P}}_{\beta\times\mathbb{D}_{<d}}$-name for a $\kappa$-complete filter on $\kappa$, where we note that ${\mathbb{P}}_{\beta\times\mathbb{D}_{<d}}={\mathbb{P}}_{<a}$.
Recall that for a $V$-generic filter $G\subseteq\mathbb{P}$, for any $a\in\mathbb{D}^{+}$, $G_{<a}$ is the induced generic for $\mathbb{P}_{<a}$. For $\alpha<\mu^{+}$ we denote by $G_{\alpha}$ the induced generic for $\mathbb{P}_{\alpha\times\mathbb{D}}$. In the next remark, we record a few easy observations, where we make repeated use of the $\kappa^{+}$-cc and the regularity of $\mu^{+}$ to analyze how subsets of $\kappa$ will appear in the generic extension.